OpenStudy (anonymous):

if f(x)=x+2 and f(g(1))=6.Which of the following could be g(x)?

OpenStudy (anonymous):

@zordoloom

OpenStudy (anonymous):

Do you not need to post the following so that we can tell you?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i need help with the solution

OpenStudy (anonymous):

you said "which of the following could be g(x)" we need the following to be able to tell you.

OpenStudy (anonymous):

okaay options 3x x+3 x-3 2x+1

OpenStudy (anonymous):

its x+3. Purely for these reasons, f(x)=x+2 and f(g(1))=6, there for if we substitute g(x) for x then we get g(1)+2=6, therefore g(1)=4, plug one into all of your potential answers and the only one that gives 4 is option 2. Voila

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Why don't you accept messages? It would be a lot easier just to directly ask me for help.

OpenStudy (anonymous):

i can ask u here also

OpenStudy (anonymous):

Touche, whatever you say.

OpenStudy (campbell_st):

g(1) has to be 4 since 4 + 2 = 6 fo g(1) = 4 for example it could be g(x) = 4x lots of choices that would give a value of 4

OpenStudy (anonymous):

in the xy plane the graph of the function is a line.If h(-1)=4 and h(5)=1 .What is the value of h(0) ?

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

@Henry.Lister

OpenStudy (campbell_st):

you need to find the equation of the line you have 2 ordered pairs (-1, 4) and (5, 1) so start by finding the slope.